Find a value of n so the linear system has an infinite number of solutions
x-y=4
-2x+2y=n

Answer by stanbon(75887)

Find a value of n so the linear system has an infinite number of solutions
x-y=4
-2x+2y=n
-------
The bottom eqution must be a multiple of the 1st equation.
Multiply the 1st equation by -2 to get:
-2x + 2y = -8
----------------
Ans: n = -8
==========================
Cheers,
Stan H.